# Enhanced sensing of Stark weak field under the influence of Aubry-André-Harper criticality

### v2

@@ Abstract
The localization-delocalization transition can be leveraged as a resource for achieving quantum-enhanced sensitivity in parameter estimation. We demonstrate that by employing different classes of localization-delocalization transition potentials one can significantly enhance the precision of parameter estimation. Specifically, we focus on the precision measurement of the Stark strength parameter encoded in the ground state of a one-dimensional fermionic lattice under the influence of Aubry-André-Harper (AAH) localization-delocalization transition. We consider the single-particle system and the system at half-filling. Our work reveals that the Quantum Fisher Information (QFI) offers a superior scaling with respect to the system size in comparison to the pure stark case, leading to a better parameter estimation. However, experimentally measuring fidelity-based QFI in a many-body system poses significant challenges. To address this, we suggest experimentally relevant operators that can be utilized to achieve precision surpassing the Heisenberg Limit (HL) or can even saturate the QFI scaling. These operators, relevant for practical experimental setups, provides a feasible pathway to harness the advantages offered by the localization-delocalization transition by exploiting two distinct localizing potentials for quantum-enhanced parameter estimation.
@@ I Introduction
The ultimate limit of precision in parameter estimation is given by the Cramér-Rao bound [1, 2, 3, 4]. For an unknown parameter, $h$, the standard deviation of parameter $\delta h$ is bounded by a relation, $\delta h\geq 1/\sqrt{MF_{Q}}$, where $M$ is the number of repetitions of the sensing protocol and $F_{Q}$ is QFI. Classically, the QFI scales with L independent qubit as $F_{Q}\sim L$, known as a standard quantum limit (SQL), whereas it can reach Heisenberg limit (HL), i.e., $F_{Q}\sim L^{2}$, for certain quantum mechanical systems [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. There are some special kind of entangled states, such as the Greenberger-Horne-Zeilinger (GHZ) state that can be exploited for achieving HL. However, GHZ-based quantum sensors are susceptible to decoherence. Recently, various kinds of quantum many-body probes have been proposed for achieving quantum-enhanced sensitivity by harnessing the different types of quantum phenomena. In Quantum-many-Body (QMB) systems, like second order, localization-delocalization and topological phase transitions has been identified as a quantum resource for parameter sensing [25, 26, 27, 7, 8, 28, 29]. Other types of many-body probes that have been reported include quantum scars [30, 31, 32, 33] and Floquet driving [12]. Generally, the precision of these probes can reach the HL. Moreover, sensing protocols with super-Heisenberg scaling have also been proposed recently on the QMB platform [34, 7, 35, 36, 37]. In majority of the quantum critical sensors proposed so far, gap closing at criticality has been identified as the key ingredient driving quantum enhanced sensing. However, recently it has been demonstrated that symmetry-breaking at criticality can also be a crucial resource for quantum sensing [38].
@@
QMB systems hosting distinct quantum phases can be used as a sensing probe. The unknown parameter to be estimated is encoded in the ground state or an excited state. These types of sensors may achieve HL at the criticality [39, 40, 41]. In order to estimate the parameter for these types of sensors we need to prepare a state and then adiabatically tune the parameter near criticality. This approach is referred as adiabatic sensing. In the critical region, the system is extremely sensitive to even a small change of the parameter. Consequently, the system has drastic response to the parameter change. As for an example, due to a self-dual symmetry, the AAH model shows localization-delocalization transition at a finite strength of the potential amplitude. This transition characterized by a quasi-periodic potential can be used as a resource for achieving quantum-enhanced parameter sensing that attains HL [42]. Recently, there has been proposals for quantum sensors using stark localization. Interesingly, one can can achieve super Heisenberg limit via Stark localization transition [43]. The QMB sensors have various applications in quantum technology [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. At the critical point or transition point the fidelity susceptibility scales with system size $L$ as $L^{2/d\nu}$, where $d$ represents the spatial dimensionality of the system, and $\nu$ is the critical exponent associated with the localization length, $\zeta$, near criticality ($h_{c}$), such that $\zeta\sim|h-h_{c}|^{-\nu}$. At the transition point the precision of the unknown parameter $h$ scales with system size as $L^{2/\nu}$ [63, 64]. Experimentally, it is difficult to measure QFI. Therefore we need to find an experimentally measurable operator. For operators, one needs to compute operator-based Fisher information (OFI) which is basically defined via signal-to-noise ratio. It can be shown that OFI is bounded by QFI [65]. In principle, the OFI saturates QFI when optimized over all possible operators.
@@
The AAH model, characterized by tight-binding model with nearest-neighbor hopping and a quasiperiodic onsite potential, undergoes a localization-delocalization phase transition when the onsite potential’s strength reaches twice the value of the hopping term [66]. The experimental evidence of this transition for AAH model has been reported in the ultracold atom setup [67, 68, 69] and in photonic crystals [70, 71, 72, 73]. On the other hand, in the tight-binding model, if a linear gradient field is induced across the lattice instead of quasiperiodic potential, the system also shows localization-delocalization transition in the limit of zero-field strength. This is the case of well-known Stark localization [74, 75, 76, 77]. In the context of the AAH model and the Stark model, the QFI exhibits distinct scaling behaviors at their respective criticalities. Specifically, at the criticality of the AAH model, the QFI scales with system size as $L^{2}$ for single particle as well as half-filled probe, which reaches the Heisenberg limit (HL). On the other hand, at the criticality of the Stark model, the QFI scales as $L^{5.9}$ for single particle probe and $L^{4.1}$ for half-filled probe, surpassing the Heisenberg limit [43]. Numerous studies have explored these critical phenomena individually in the presence of weak interactions [78, 79]. Recent studies have investigated the characteristic properties of fundamental physical quantities, such as localization length, inverse participation ratio, and energy gap, by examining systems in presence of these two different types of potentials. These studies reveal that the scaling exponents associated with the localization length, energy gap, and other related quantities are markedly different from the pure cases [80]. In this work, we investigate the advantage of interplay between these two potentials in the quantum sensing of stark weak field while initiating the system at the AAH criticality. Our results demonstrate that the precision of the parameter encoded in the Stark field potential can be significantly enhanced when measuring the parameter $h$ in the presence of the AAH criticality. This suggests a novel approach to improve parameter estimation in quantum systems by leveraging the critical properties of different classes of quantum phase transitions, in general.
@@
In the following, section I present the introduction. Section II presents the fidelity-based definition of QFI and OFI. The system is introduced in section III. Section IV demonstrates the behavior of OFIs and their scaling with system size for a single-particle probe. Section V covers the QFI, OFIs, and their scaling for a many-body half-filled probe. Finally, the discussion is provided in section VI.
@@
where ${\mathcal{F}}_{Q}=|\langle\psi(h)|\psi(h+\delta h)\rangle|$ is the distance between two nearby states $|\psi(h)\rangle$ and $|\psi(h+\delta h)\rangle$. The QFI is related to fidelity susceptibility as ${\mathcal{F}}=4\chi_{Q}$ [8]. The estimation of an unknown parameter in experiment requires detecting average of an operator $\hat{O}$, $\langle\hat{O}\rangle=\text{Tr}[\hat{O}\rho(h)]$ in the function of $h$. where $\rho(h)$ is the quantum state of the sensor. The sensitivity of $h$ is quantified by the error propagation formula, which is essentially the signal-to-noise ratio, defined as
@@
where $\text{Var}(\hat{O})$ is defined as, $\text{Var}(\hat{O})=\langle{\hat{O}}^{2}\rangle-\langle{\hat{O}}\rangle^{2}$. The quantum Cramér-Rao bound [3, 4, 65] provides the bound on the uncertainty for any observable estimation: $F_{O}(h,\hat{O})\leq F_{Q}(h)$.
@@ III System
We consider a one-dimensional lattice with nearest-neighbor hopping in the presence of a quasi-periodically modulated potential and a linear gradient field, the so-called stark field. The generic form of Hamiltonian is $\hat{H}=\hat{H_{0}}+h\hat{H_{2}}$. The explicit form of the Hamiltonian is given by,
$\hat{H_{0}}$ $=-\sum_{i}^{L-1}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\mathrm{h.c.})+V\sum_{i}^{L-1}\cos[2\pi(i\omega+\phi)]\hat{c}^{\dagger}_{i}\hat{c}_{i},$
$\hat{H_{1}}$ $=h\sum_{i}^{L-1}i\hat{c}^{\dagger}_{i}\hat{c}_{i}.$ (3)
@@
where $\hat{c}_{i}^{\dagger}(\hat{c}_{i})$ is creation (annihilation) operator and $L$ denotes the system size. The first term in $\hat{H}_{0}$ is the kinetic energy term. The lattice is subjected to two different kinds of onsite potentials. The second term in $\hat{H}_{0}$ corresponds to a quasiperiodic onsite potential and $H_{1}$ has a contribution from the onsite potential energy due to the linear gradient field with strength $h$. Here $\omega$ is taken as $\omega=F_{n}/F_{n+1}$, where $F_{n}$ is the $n^{\text{th}}$ Fibonacci number. For $n\to\infty$ limit, $\omega=\lim_{n\to\infty}F_{n}/F_{n+1}\to(\sqrt{5}-1)/2$, which is golden ratio. In the presence of the quasiperiodic potential, appropriate scaling emerges for the system sizes belonging to the Fibonacci series [80]. $\phi$ is a phase which is chosen randomly from a uniform distribution in [0,1]. We consider open boundary condition (OBC). The Hamiltonian ($\hat{H}$) reduces to pure stark case for $V=0$. Then the system is localized in the zero field limit, i.e., $h\to 0$. At this transition point the fidelity susceptibility scales with system size as $\chi_{Q}\sim L^{2/\nu}$ with $\nu=0.33$ for single particle ground state [43]. Now, when the system is additionally at the criticality of AAH, i.e., $V=2$, the corresponding Hamiltonian is under the influence of two distinct onsite potentials. The scaling exponent associated with the localization-delocalization is reported as $\nu=0.29$ [80].
@@
Figure 1: Single party OFI: (a) presents OFI, $F_{cdw}$, for observable $\hat{O}_{cdw}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (b) presents OFI, $F_{H_{2}}$ for observable $\hat{O}_{H_{2}}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (c) illustrates the scaling of OFI corresponding to the operator $\hat{O}_{cdw}$, $F_{cdw}$ (circles) and $F_{H_{2}}$ (stars), respectively. The straight lines are the best fits. Whereas $F_{cdw}$ scales as $F_{cdw}\sim L^{6.2}$, $F_{H_{2}}$ saturates the scaling of the QFI, i.e., $F_{H_{2}}\sim L^{6.7}$.
IV Single particle system as probe
We first consider the single-particle scenario. We aim to precisely measure $h$ by setting the AAH potential at criticality. All eigenfunctions undergo delocalized to localized phase transition for this system. In the thermodynamic limit ($L\to\infty$) the transition point takes place at $h\to 0$ [43]. To address whether the sensitivity of $h$ can be enhanced using the criticality of the AAH model, we investigate and calculate the OFI in detail.
The scaling of QFI with finite system sizes at the extended region is $F_{Q}\sim L^{\beta}$ with $\beta=6.7$ which is shown in the reference [80]. In this section, we have shown that the experimentally measurable operator can also achieve super-Heisenberg scaling for measurement precision of the parameter $h$.
Figure 2: Half-filled case QFI: (a1) QFI for half filled case with respect to $h$. We have taken system size $L(n_{f})$, where $L$ is system size and $n_{f}$ is the number of filling. This figure shows that in the flat region, the QFI is increasing with system sizes, $L$ = 55 (28) (blue dash-dash), 89 (44) (red dot-dash), 233 (116) (black dot-dot-dash), 377 (188) (green dash-dash-dot). After a certain value of $h$, the QFI decays with certain initial fluctuations and again it shows steady behavior. (a2) In this plot we have shown the scaling of QFI with system size $L(n_{f})$. The circle point is numerical value of QFI for $h=10^{-9}$ for $L$ = 21 (10), 55 (28), 89 (44), 233 (116), 377 (188), 987 (493) and the dotted line is the best fitting with fitting function $F_{Q}(h=10^{-9})\sim L^{6.6}$.
Figure 3: Half-filled case OFI: (b1) The figure illustrates the behavior of the OFI for system size $L(n_{f})$ ($n_{f}$ number of filling) with respect to $h$. In the flat region, the OFI increases with system size. However, after reaching a certain value of $h$, the OFI begins to decay with fluctuations and eventually stabilizes into a steady behavior. We have taken $L$ = 55 (blue dash-dash), 89 (red dot-dash), 233 (black dot-dot-dash), 377 ( green dash-dash-dot). (b2) In this plot, we show the scaling of OFI with system size. The violet circular points represent the numerical values of OFI for $h=10^{-9}$ with $L$ = 21 (10), 55 (28), 89 (44), 233 (116), 377 (188), 987 (493), while the violet dotted line represents the best fit. We obtain $F_{H_{2}}(h=10^{-9})\sim L^{6.6}$.
IV.1 Operator Fisher Information
Experimentally accessing the QFI poses significant challenges. So, we need to identify an observable that is both experimentally relevant and provides a significant quantum advantage in parameter estimation. We focus on the operator $\hat{O}_{cdw}=\sum_{i}(-1)^{i}\hat{c}_{i}^{\dagger}\hat{c}_{i}$, which measures the occupation imbalance between even and odd sites. This operator reveals the charge-density-wave (CDW) order in a quantum state and can be directly measured in optical lattice setups with ultracold atoms [67, 68, 69]. The computation of OFI is followed by the Eq. (2), which entails the average of $\langle O_{cdw}\rangle$ and $\langle\hat{O}_{cdw}\rangle^{2}$, where, $\langle\hat{O}_{cdw}\rangle^{2}=\sum_{i,j}(-1)^{i+j}\hat{c}_{i}^{\dagger}\hat{c}_{i}\hat{c}_{j}^{\dagger}\hat{c}_{j}$.

### v3

# Enhanced sensing of a weak Stark field under the influence of Aubry-André-Harper criticality

@@ Abstract
The localization transition can be exploited as a resource for achieving quantum-enhanced sensitivity in parameter estimation. We demonstrate that by employing different classes of localization inducing potentials, one can significantly enhance the precision of parameter estimation. Specifically, we focus on the precision measurement of the Stark strength parameter encoded in the low- and high-energy eigenstates of a one-dimensional fermionic lattice under the influence of Aubry-André-Harper localization-delocalization transition. For the ground state, we consider the single-particle system, in addition to the system at half filling. Our work reveals that Quantum Fisher Information (QFI) offers superior scaling with respect to the system size compared to the pure stark case, leading to a better parameter estimation. However, experimental measurement of the QFI based on fidelity in a multibody system is a significant challenge. To address this, we suggest experimentally relevant operators that can be utilized to achieve precision surpassing the Heisenberg Limit (HL) or can even saturate the QFI scaling. These operators, relevant for practical experimental setups, provide a feasible pathway to harness the advantages offered by the localization-delocalization transition by exploiting two distinct localizing potentials for quantum-enhanced parameter estimation. In addition, to investigate the robustness of the proposed quantum sensors, we add noise in the form of thermal fluctuations, and demonstrate that the Fisher information of the thermal states can saturate to HL. In principle, this work demonstrates how introducing additional quantum-criticality-inducing control parameters could be utilized to enhance quantum sensitivity in the estimation of an unknown target parameter, and hence promises applications in wider contexts, e.g., in hybrid systems involving other classes of localization inducing potential, such as Anderson localization, and even in systems involving other types of quantum criticalities, such as second-order and topological phase transitions.
@@ I Introduction
The ultimate limit of precision in parameter estimation is given by the Cramér-Rao bound [1, 2, 3, 4]. For an unknown parameter, $h$, the standard deviation of parameter $\delta h$ is bounded by the relation, $\delta h\geq 1/\sqrt{MF_{Q}}$, where $M$ is the number of repetitions of the sensing protocol and $F_{Q}$ is the quantum Fisher information (QFI). Classically, the QFI scales with $L$ independent qubit as $F_{Q}\sim L$, known as a standard quantum limit (SQL), whereas it can reach the Heisenberg limit (HL), i.e., $F_{Q}\sim L^{2}$, for certain quantum mechanical systems [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. There are some special kind of entangled states, such as the Greenberger-Horne-Zeilinger (GHZ) state that can be exploited for achieving the HL. However, GHZ-based quantum sensors are susceptible to decoherence. Recently, various kinds of quantum many-body probes have been proposed for achieving quantum-enhanced sensitivity by harnessing the different types of quantum phenomena. Quantum many-Body (QMB) systems, such as second-order, localization-delocalization, and topological phase transitions have been identified as a quantum resource for parameter sensing [27, 28, 29, 7, 8, 30, 31]. Other types of many-body probes that have been reported include quantum scars [32, 33, 34, 35] and Floquet driving [12]. Generally, the precision of these probes can reach the HL. Moreover, sensing protocols with super-Heisenberg scaling have also been proposed recently on the QMB platform [36, 7, 37, 38, 39]. There is some nuance in the behaviors that constitute a resource in the critical parameter-estimation sensing: it is the confluence of diverging susceptibility, enhanced responses to small perturbations, and the generation of long-range correlations, symmetry breaking, and, in some circumstances, gap closing, which can lead to an enhancement [26, 40, 41].
@@
QMB systems hosting distinct quantum phases can be used as a sensing probe. One may consider estimating a parameter that is encoded in the ground state or in an excited state of a system. These types of sensors may achieve the HL at the criticality [42, 43, 44]. In order to estimate an unknown parameter governing the quantum phase transition in these types of sensors, we need to prepare a state and then adiabatically tune the parameter to be estimated near criticality. This approach is referred as adiabatic sensing. In the critical region, the system becomes sensitive to small change to the parameter, resulting in a divergent quantum Fisher information. As an example, due to a self-dual symmetry, the Aubry-André-Harper (AAH) model shows localization-delocalization transition at a finite strength of the potential amplitude. This transition characterized by a quasiperiodic potential can be used as a resource for achieving quantum-enhanced parameter sensing that attains the HL [45]. Recently, there have been proposals for quantum sensors using Stark localization. Interestingly, one can achieve the super Heisenberg limit via Stark localization transition [46]. QMB sensors have various applications in quantum technology [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. At the critical point or transition point the fidelity susceptibility scales with system size $L$ as $L^{2/d\nu}$, where $d$ represents the spatial dimensionality of the system, and $\nu$ is the critical exponent associated with the localization length, $\zeta$, near criticality ($h_{c}$), such that $\zeta\sim|h-h_{c}|^{-\nu}$. At the transition point, the precision of the unknown parameter $h$ scales with system size as $L^{2/\nu}$ [66, 67]. Implementing the optimal measurement that saturates the QFI is often challenging in practice. Therefore, it is necessary to identify an operator that is both theoretically meaningful and experimentally accessible. For operators, one needs to compute operator-based Fisher information (OFI), which is basically defined via signal-to-noise ratio. It can be shown that OFI is bounded by QFI [68]. In principle, the OFI saturates QFI when optimized over all possible operators.
@@
The AAH model, characterized as a tight-binding model with nearest-neighbor hopping and a quasiperiodic onsite potential, undergoes a localization-delocalization phase transition when the onsite potential’s strength reaches twice the value of the hopping term [69, 70, 71]. The experimental evidence of this transition for the AAH model has been reported in ultracold atom setup [72, 73, 74] and in photonic crystals [75, 76, 77, 78]. On the other hand, in the tight-binding model, if a linear gradient field is induced across the lattice instead of quasiperiodic potential, the system also shows a localization-delocalization transition in the limit of zero field strength. This is the case of well-known Stark localization [79, 80, 81, 82]. In the context of the AAH model and the Stark model, the QFI exhibits distinct scaling behaviors at their respective criticalities. Specifically, at the criticality of the AAH model, the QFI scales with system size as $L^{2}$ for single particle as well as half-filled probe, which reaches the HL. On the other hand, at the criticality of the Stark model, the QFI scales as $L^{5.9}$ for single particle probe and $L^{4.1}$ for a half-filled probe, surpassing the Heisenberg limit [46]. Numerous studies have explored these critical phenomena individually in the presence of weak interactions [83, 84]. Recent studies have investigated the characteristic properties of fundamental physical quantities, such as localization length, inverse participation ratio, and energy gap, by examining systems in the presence of these two different types of potentials. These studies reveal that the scaling exponents associated with the localization length, energy gap, and other related quantities are markedly different from the pure cases [85]. In this work, we investigate the advantage of interplay between these two potentials in the quantum sensing of a weak Stark field while initiating the system at the AAH criticality. Our results demonstrate that the precision of the parameter encoded in the Stark field potential can be significantly enhanced when measuring the parameter $h$ in the presence of the AAH criticality. This suggests a novel approach to improve parameter estimation in quantum systems by leveraging the critical properties of different classes of quantum phase transitions, in general.
@@
The introduction is presented in Sec. I. Section II presents the fidelity-based definitions of QFI and OFI. The system is introduced in Sec. III. Sections IV and V examine the behavior of QFI and OFI, as well as their scaling with system size, for a single-particle probe in the ground and mid-spectrum states, respectively. Section VI explores the behavior of the QFI and OFI for the thermal states. Section VII investigates QFI, OFIs, and their scaling for a many-body half-filled probe. Finally, Sec. VIII provides a Conclusion.
@@
where ${\mathcal{F}}_{Q}=|\langle\psi(h)|\psi(h+\delta h)\rangle|$ is the distance between two nearby states $|\psi(h)\rangle$ and $|\psi(h+\delta h)\rangle$. The QFI is related to fidelity susceptibility as ${\mathcal{F}}=4\chi_{Q}$ [8]. Estimating an unknown parameter requires measuring an observable whose expectation value depends on that parameter. For an operator $\hat{O}$, $\langle\hat{O}\rangle=\text{Tr}[\hat{O}\rho(h)]$ is the function of $h$. where $\rho(h)$ is the quantum state of the sensor. The sensitivity of $h$ is quantified by the error propagation formula, which is essentially the signal-to-noise ratio, defined as
@@
[…]
@@ III System
We consider a one-dimensional lattice with nearest neighbor hopping in the presence of a quasiperiodically modulated potential (Aubry–André) and a linear gradient (Stark) field, referred to as the Aubry–André–Stark (AAS) model. The generic form of Hamiltonian is $\hat{H}=\hat{H_{1}}+h\hat{H_{2}}$. The explicit form of the Hamiltonian is given by,
$\hat{H_{1}}$ $=-\sum_{i}^{L-1}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\mathrm{H.c.})+V\sum_{i}^{L-1}\cos[2\pi(i\omega+\phi)]\hat{c}^{\dagger}_{i}\hat{c}_{i},$
$\hat{H_{2}}$ $=h\sum_{i}^{L-1}i\hat{c}^{\dagger}_{i}\hat{c}_{i}.$ (3)
@@
where $\hat{c}_{i}^{\dagger}(\hat{c}_{i})$ is the creation (annihilation) operator and $L$ denotes the system size. The first term in $\hat{H}_{1}$ is the kinetic energy term. The lattice is subjected to two different kinds of onsite potentials. The second term in $\hat{H}_{1}$ corresponds to a quasiperiodic onsite potential, and $\hat{H}_{2}$ has a contribution from the onsite potential energy due to the linear gradient field with strength $h$. Here $\omega$ is taken as $\omega=F_{n}/F_{n+1}$, where $F_{n}$ is the $n{\text{th}}$ Fibonacci number. For the $n\to\infty$ limit, $\omega=\lim_{n\to\infty}F_{n}/F_{n+1}\to(\sqrt{5}-1)/2$, which is the golden ratio. In the presence of the quasiperiodic potential, appropriate scaling emerges for the system sizes belonging to the Fibonacci series [85]. $\phi$ is a phase which is chosen randomly from a uniform distribution in [0,1]. We consider an open boundary condition (OBC).
@@
Figure 1: Single party OFI. (a) presents OFI, $F_{cdw}$, for observable $\hat{O}_{cdw}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (b) presents OFI, $F_{H_{2}}$ for observable $\hat{O}_{H_{2}}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (c) illustrates the scaling of OFI corresponding to the operator $\hat{O}_{cdw}$, $F_{cdw}$ (circles) and the operator $\hat{O}_{H_{2}}$, $F_{H_{2}}$ (stars) for L = 21, 34, 55, 89, 144, 233, 377, 610, 987, 1597 . The straight lines are the best fits. Whereas $F_{cdw}$ scales as $F_{cdw}\sim L^{6.2}$, $F_{H_{2}}$ saturates the scaling of the QFI, i.e., $F_{H_{2}}\sim L^{6.7}$. A configuration averaging over $\phi$ is performed with 8000 random samples.
IV Single-particle ground state as probe
Quantum Fisher information. We first consider the single-particle scenario. The Hamiltonian ($\hat{H}$) reduces to the pure Stark case for $V=0$ (i.e., in the absence of a quasiperiodic potential). In this limit, the system becomes localized as the field strength approaches zero, $h\to 0$. At this transition point, the fidelity susceptibility scales with system size as $\chi_{Q}\sim L^{2/\nu}$ with $\nu=0.33$ for the single-particle ground state, and the QFI exhibits the scaling $F_{Q}\sim L^{6}$ [46]. When the system is additionally tuned to the critical point of the AAH model, i.e., $V=2$, the Hamiltonian experiences the combined effect of two distinct onsite potentials. In this case, it is known from our previous work [85] that the scaling exponent associated with the localization–delocalization transition turns out to be $\nu=0.29$, and the QFI in the delocalized regime scales as $F_{Q}\sim L^{\beta}$ with $\beta=6.7$.
[…]
Observable Fisher information. Despite the fact that the QFI continues to provide the ultimate bound on estimation precision, independent of measurement strategy, experimentally accessing the QFI poses significant challenges. In order to overcome this practical constraint, we need to identify an observable that is both experimentally relevant and provides a significant quantum advantage in parameter estimation, which forms the central theme of this work.